# Transverse spin effects and light-quark dipole moments at lepton colliders

### v1

@@ Abstract
Introduction.—Measuring electroweak dipole moments of particles is crucial for testing the Standard Model (SM) and probing New Physics (NP). Electroweak dipole moments are related to $CP$ violation, needed to generate the observed baryon asymmetry of the universe [1], and can be systematically parameterized by dimension-six operators within the framework of the SM Effective Field Theory (SMEFT) [2, 3]. These operators introduce NP interactions at the scale $\Lambda$, which are characterized by fermion helicity-flip effects in high-energy scattering and could be the origin of certain observed anomalies at the Large Hadron Collider (LHC). One example is the observed Lam-Tung relation breaking in the Drell-Yan process [4, 5, 6, 7, 8, 9, 10, 11]. However, because of this property, in the current global analyses of SMEFT operators that only use unpolarized production rates [12, 13, 14, 15, 16, 17, 18], to which the contributions from light-fermion dipole operators start only quadratically at $\mathcal{O}(1/\Lambda^{4})$ or are significantly suppressed by the small fermion mass(es) in their interference with the SM amplitudes at $\mathcal{O}(1/\Lambda^{2})$, the corresponding Wilson coefficients are poorly constrained.
@@
It is the purpose of this Letter to examine the possibility of constraining light-quark dipole couplings at a future lepton-lepton collider. This can be viewed as a crossed process of dihadron production in $ep$ collisions, where the initial-state proton is flipped to the final state, with a much cleaner background. Furthermore, the experimental observable is no longer limited to the proton-dihadron final state. As we will show, a combined analysis of the associated production of a $(\pi^{+}\pi^{-})$ pair with another hadron $h^{\prime}$, e.g., $h^{\prime}=\pi^{\pm}$, $K^{\pm}$, $p$ or $\bar{p}$, enables us to disentangle the dipole couplings of up and down quarks.
Observable and factorization.—We consider the inclusive associated production of a dihadron-hadron pair in $e^{-}e^{+}$ collisions, in the center-of-mass (c.m.) frame,
@@
where $y=(1-\cos\theta)/2$ trades for the polar angle $\theta$ of the dihadron, and $(z,\bar{z})=2(P_{h}^{0},p^{\prime 0})/\sqrt{s}$ are the fractional energies. Then we can write the cross section of Eq. (1) in a factorized form,
@@
The factorization formula in Eq. (3) separates the physical stages at different scales, as depicted in Fig. 2. At the leading order (LO), the hard scattering produces a quark and antiquark, $e^{-}_{\lambda_{1}}e^{+}_{\lambda_{2}}\to q_{\lambda}\bar{q}_{\bar{\lambda}}$, whose helicity amplitude $\mathcal{M}^{\lambda_{1}\lambda_{2}}_{\lambda\bar{\lambda}}$ determines the hard coefficient $C_{q}$, which controls the unpolarized production rate, and the transverse spin $\bm{s}_{T,q}=(s^{x}_{q},s^{y}_{q})$ of the quark through the (unnormalized) density matrix,
@@
where $\lambda$’s are the helicity indices with repeated ones summed over, $\sigma^{i}=(\sigma^{x},\sigma^{y},\sigma^{z})$ are the Pauli matrices, and $s_{q}^{i}=(s_{q}^{x},s_{q}^{y},s_{q}^{z})$ is the quark spin vector defined in the dihadron frame, cf. Eq. (2). In the case when the initial-state leptons are polarized, we shall average their helicities in Eq. (4) with proper density matrices.
@@
Dipole interaction and transverse spin.—As shown in Eq. (4), the transverse spin $\bm{s}_{T,q}$ of our interest is given by the off-diagonal element of the density matrix,
@@
where $q_{L}=(u_{L},d_{L})^{T}$ is the first-generation left-handed quark doublet, $u_{R}$ and $d_{R}$ are the right-handed singlets, and $H$ and $\widetilde{H}=i\sigma^{2}H^{*}$ are the Higgs doublets, with $v=246~{}{\rm GeV}$ being the vacuum expectation value. The $B_{\mu\nu}$ and $W_{\mu\nu}^{a}$ are the field strength tensors of the gauge groups U$(1)_{Y}$ and SU$(2)_{L}$, respectively, and $g_{1}$ and $g_{2}$ are the corresponding gauge couplings. The dimensionless Wilson coefficients $\Gamma_{B}^{u,d}$ and $\Gamma_{W}^{u,d}$ are the dipole interaction couplings. For convenience, we redefine the $u$ and $d$ quark dipole couplings in terms of the gauge boson mass eigenstates as $\Gamma_{\gamma}^{u,d}=\mp\Gamma_{W}^{u,d}-\Gamma_{B}^{u,d}$ and $\Gamma_{Z}^{u,d}=\mp c_{W}^{2}\Gamma_{W}^{u,d}+s_{W}^{2}\Gamma_{B}^{u,d}$, where the $\mp$” corresponds to $u$ or $d$, respectively, and $s_{W}\equiv\sin\theta_{W}$ and $c_{W}\equiv\cos\theta_{W}$, with $\theta_{W}$ being the weak mixing angle.
@@
with a similar relation for antiquark. At the LO, the two coefficients $w^{q}_{\gamma,Z}$ are real functions of the hard scattering kinematics, [^2] We neglect the imaginary part of $w_{\gamma,Z}^{q}$ from the $Z$-width, as it is suppressed. so the transverse spins are simply given by
@@
By measuring the quark transverse spin through the $\phi_{R}$ modulation in Eq. (5), one gains sensitivity to the dipole couplings—both real and imaginary parts—at $\mathcal{O}(1/\Lambda^{2})$. This greatly enhances the sensitivity with only unpolarized rates, for which one integrates over $\phi_{R}$ in Eq. (3) and only accesses the $C_{q}\propto|\mathcal{M}^{\lambda_{1}\lambda_{2}}_{\lambda\bar{\lambda}}|^{2}$, where the helicity-flipping dipole interactions have to appear in pairs so contribute only quadratically at $\mathcal{O}(1/\Lambda^{4})$.
@@
Azimuthal asymmetry and flavor constraint.—There is currently only the $(h_{1}h_{2})=(\pi^{+}\pi^{-})$ DiFFs at our disposal, taken from the global-fit results in Refs. [39, 40, 41]. These satisfy the following flavor relations due to isospin and charge conjugation symmetries:
@@
[…]
with $i=x$ or $y$. Since the differences between $(C_{q},s_{q}^{i})$ and $(C_{\bar{q}},s_{\bar{q}}^{i})$, as derived from Eq. (9), are due to the axial coupling $g_{A}^{q}$ of the $Z$-$q$-${\bar{q}}$ interaction, which breaks charge symmetry and induces a forward-backward asymmetry, they vanish under the integration of $y$, and hence $\left\langle C_{q}\right\rangle=\left\langle C_{\bar{q}}\right\rangle$ and $\left\langle S_{q}^{i}\right\rangle=\left\langle S_{\bar{q}}^{i}\right\rangle$. This, together with Eq. (10), allows to write Eq. (3) in a concise form,
@@
That leveraging the associatively produced hadron $h^{\prime}$ in a multi-channel analysis enables one to disentangle the quark flavors of the dipole couplings is the most important observation in this Letter. Without measuring $h^{\prime}$, which amounts to removing the $\bar{z}$ dependence and setting the FF to unity in Eq. (3), the $B^{i}$ vanishes identically in Eq. (14). To restore the transverse spin effects in this case requires not to integrate over $y$ in Eq. (11), and to utilize the forward-backward asymmetry associated with the charge symmetry breaking effects. However, such observable still constrains only a linear combination of the up and down quark dipole couplings, similarly to Refs. [20, 22], but not the individual ones.
Numerical results and discussion.—To simplify the numerical analysis, we further integrate over $(z,\bar{z},M_{h})$ in Eq. (12), resulting in a single-differential $\phi_{R}$ distribution,
$\frac{2\pi}{\sigma^{h^{\prime}}}\frac{d\sigma^{h^{\prime}}}{d\phi_{R}}=1-a_{R}^{h^{\prime}}\sin\phi_{R}+a_{I}^{h^{\prime}}\cos\phi_{R}+\order{\Lambda^{-4}},$ (15)
@@
As discussed, in order to separate the photon and $Z$ dipole couplings, we propose to carry out the measurements both at low-energy lepton colliders, such as the Super Tau-Charm Facility (STCF) [51], BELLE [52], and high-energy $Z$ factories, such as the Circular Electron-Position Collider (CEPC) [53, 54], International Linear Collider (ILC) [55], and the Future Circular Collider (FCC-ee) [56]. As a benchmark, we choose $\sqrt{s}=10\,{\rm GeV}$ with a polarized electron beam, with $\lambda_{\ell}=0.8$, to enhance the signal from the $\gamma\gamma$ channel to probe $\Gamma^{u,d}_{\gamma}$, and $\sqrt{s}=91\,{\rm GeV}$, at the $Z$-pole, with unpolarized lepton beams to probe the $\Gamma^{u,d}_{Z}$ from the $ZZ$ channel, both with an integrated luminosity of $\mathcal{L}=1000~{}{\rm fb}^{-1}$. At either energy, we choose the range of $(z,\bar{z},M_{h})$ to agree with the global-fit analyses [43, 46, 39, 40, 41] and satisfy the factorization condition in Eq. (3), i.e., $z\in(0.19,0.99)$, $\bar{z}\in(0.2,0.8)$, and $M_{h}\in(0.28,2.05)\,{\rm GeV}$.
@@
In Fig. 3, we present the expected constraints on the light-quark dipole couplings $(\Gamma_{i}^{u},\Gamma_{i}^{d})$ for (a) $i=\gamma$ and (b) $i=Z$, at a 68% confidence level, separately from the $\pi^{\pm}$ (red bands), $K^{\pm}$ (blue bands), and $p/\bar{p}$ (green bands) associated production channels. The real and imaginary parts are constrained in the same way. As expected, the linear dependence of $A_{LR,UD}^{h^{\prime}}$ on the dipole couplings leads to straight-band confidence regions; the $K^{\pm}$ channel only probes the $\Gamma_{\gamma,Z}^{u}$; and the $\pi^{\pm}$ and $p/\bar{p}$ give opposite signs for the slopes of the constraint bands. These multiple channels significantly constrain the dipole couplings, yielding a typical upper limit of $\mathcal{O}(10^{-2})$ for $\Gamma_{\gamma}^{u,d}$ and $\mathcal{O}(10^{-3})$ for $\Gamma_{Z}^{u,d}$, simultaneously for their real and imaginary parts. This complements the constraints from an $ep$ collider, especially when comparing with their flat directions of the constraining contour in the $(\Gamma_{i}^{u},\Gamma_{i}^{d})$ space [20, 22].
@@
While the dipole couplings $\Gamma_{\gamma,Z}^{u,d}$ can also be constrained by the Drell-Yan process, Higgs boson production via vector boson fusion, diboson production, associated production of weak and Higgs bosons at the LHC, low-energy data, and electroweak precision observables, the dipole couplings are only constrained quadratically at $\mathcal{O}(1/\Lambda^{4})$ when the transverse spin information associated with the initial or final state of the hard scattering processes is not considered. As a result, the sensitivity of those methods, among which Drell-Yan data provide the strongest constraint, is weaker than our results by one to two orders of magnitude, even with only one operator considered at a time [12, 13, 14, 15]. Furthermore, these methods lack the sensitivity to distinguish between the real and imaginary parts of dipole couplings and cannot separate them from other SMEFT operators. This highlights the advantage of the transverse spin observable proposed in this Letter, which depends linearly and exclusively on the SMEFT dipole operators.
@@
Conclusion.—In this Letter, we introduced a novel approach to probe the light-quark dipole interactions through the azimuthal asymmetry between a collinear pair of hadrons produced in association with a well-separated third hadron at $e^{-}e^{+}$ colliders. This asymmetry results from the transverse spin of a final-state light quark, and is exclusively sensitive to the interference of the dipole and the SM interactions at the leading power of $\mathcal{O}(1/\Lambda^{2})$, unaffected by other new physics interactions in the hard scattering. This approach can simultaneously constrain the real and imaginary parts of the dipole couplings, provide a new way to probe $CP$ violation effects at $e^{-}e^{+}$ colliders, and outperform conventional methods proposed at the LHC and other observables at the $e^{-}e^{+}$ colliders, by one to two orders of magnitude. By leveraging a hadron $h^{\prime}$ that is produced in association with the dihadron pair and combining all possible channels of $h^{\prime}$ ($\pi^{\pm}$, $K^{\pm}$, and $p/\bar{p}$), we demonstrated that the $e^{-}e^{+}$ collider can resolve the flavor degeneracy at $ep$ collider, giving individual constraints on the up and down quark dipole couplings.

### v2

@@ Abstract
I Introduction
Measuring electroweak dipole moments of particles is crucial for testing the Standard Model (SM) and probing New Physics (NP). Electroweak dipole moments are related to $CP$ violation, needed to generate the observed baryon asymmetry of the Universe [1], and can be systematically parametrized by dimension-six operators within the framework of the SM effective field theory (SMEFT) [2, 3]. These operators introduce NP interactions at the scale $\Lambda$, which are characterized by fermion helicity-flip effects in high-energy scattering and could be the origin of certain observed anomalies at the Large Hadron Collider (LHC). One example is the observed Lam-Tung relation breaking in the Drell-Yan process [4, 5, 6, 7, 8, 9, 10, 11]. However, because of this property, in the current global analyses of SMEFT operators that only use unpolarized production rates [12, 13, 14, 15, 16, 17, 18], to which the contributions from light-fermion dipole operators start only quadratically at $\mathcal{O}(1/\Lambda^{4})$ or are significantly suppressed by the small fermion mass(es) in their interference with the SM amplitudes at $\mathcal{O}(1/\Lambda^{2})$, the corresponding Wilson coefficients are poorly constrained.
@@ I Introduction
It is the purpose of this paper to examine the possibility of constraining light-quark dipole couplings at a future lepton-lepton collider. This can be viewed as a crossed process of dihadron production in $ep$ collisions, where the initial-state proton is flipped to the final state, with a much cleaner background. Furthermore, the experimental observable is no longer limited to the proton-dihadron final state. As we will show, a combined analysis of the associated production of a $(\pi^{+}\pi^{-})$ pair with another hadron $h^{\prime}$, e.g., $h^{\prime}=\pi^{\pm}$, $K^{\pm}$, $p$ or $\bar{p}$, enables us to disentangle the dipole couplings of up and down quarks.
II Observable and factorization
We consider the inclusive associated production of a dihadron-hadron pair in $e^{-}e^{+}$ collisions, in the center-of-mass (c.m.) frame,
@@ II Observable and factorization
where $y=(1-\cos\theta)/2$ trades for the polar angle $\theta$ of the dihadron, and $(z,\bar{z})=2(P_{h}^{0},p^{\prime 0})/\sqrt{s}$ are the fractional energies. Then we can write the cross section of Eq. (1) in a factorized form [31, 38],
@@
The factorization formula in Eq. (3) separates the physical stages at different scales, as depicted in Fig. 2. At the leading order (LO) [^2] Higher-order QCD corrections are not expected to alter the conclusions of our analysis, analogous to the cancellation effect observed in the spin asymmetry ratio for the polarized Drell-Yan process [39]. While dihadron pairs can also arise from polarized gluons, such contributions can only generate $\cos 2\phi_{R}$ and/or $\sin 2\phi_{R}$ modulations. Consequently, the $\cos\phi_{R}$ and $\sin\phi_{R}$ asymmetries represent robust signatures of transversely polarized quarks—an attribute unaffected by gluon radiation due to chiral symmetry. the hard scattering produces a quark and antiquark, $e^{-}_{\lambda_{1}}e^{+}_{\lambda_{2}}\to q_{\lambda}\bar{q}_{\bar{\lambda}}$, whose helicity amplitude $\mathcal{M}^{\lambda_{1}\lambda_{2}}_{\lambda\bar{\lambda}}$ determines the hard coefficient $C_{q}$, which controls the unpolarized production rate, and the transverse spin $\bm{s}_{T,q}=(s^{x}_{q},s^{y}_{q})$ of the quark through the (unnormalized) density matrix,
@@
where the $\lambda$’s are the helicity indices with repeated ones summed over, $\sigma^{i}=(\sigma^{x},\sigma^{y},\sigma^{z})$ are the Pauli matrices, and $s_{q}^{i}=(s_{q}^{x},s_{q}^{y},s_{q}^{z})$ is the quark spin vector defined in the dihadron frame, cf. Eq. (2). In the case when the initial-state leptons are polarized, we shall average their helicities in Eq. (4) with proper density matrices.
@@
III Dipole interaction and transverse spin
As shown in Eq. (4), the transverse spin $\bm{s}_{T,q}$ of our interest is given by the off-diagonal element of the density matrix,
@@ III Dipole interaction and transverse spin
[…]
@@
with a similar relation for the antiquark. At the LO, the two coefficients $w^{q}_{\gamma,Z}$ are real functions of the hard scattering kinematics, [^3] We neglect the imaginary part of $w_{\gamma,Z}^{q}$ from the $Z$-width, as it is suppressed. so the transverse spins are simply given by
@@
By measuring the quark transverse spin through the $\phi_{R}$ modulation in Eq. (5), one gains sensitivity to the dipole couplings—both the real and imaginary parts—at $\mathcal{O}(1/\Lambda^{2})$. This greatly enhances the sensitivity with only unpolarized rates, for which one integrates over $\phi_{R}$ in Eq. (3) and only accesses the $C_{q}\propto|\mathcal{M}^{\lambda_{1}\lambda_{2}}_{\lambda\bar{\lambda}}|^{2}$, where the helicity-flipping dipole interactions have to appear in pairs so contribute only quadratically at $\mathcal{O}(1/\Lambda^{4})$.
@@
IV Azimuthal asymmetry and flavor constraint
There is currently only the $(h_{1}h_{2})=(\pi^{+}\pi^{-})$ DiFFs at our disposal, taken from the global-fit results in Refs. [41, 42, 43]. These satisfy the following flavor relations due to isospin and charge conjugation symmetries:
@@ IV Azimuthal asymmetry and flavor constraint
$\left\langle C_{q}\right\rangle=\int_{0}^{1}dy\,C_{q}(y),\quad\left\langle S_{q}^{i}\right\rangle=\int_{0}^{1}dy\,C_{q}(y)s_{q}^{i}(y),$ (11)
with $i=x$ or $y$. Since the differences between $(C_{q},s_{q}^{i})$ and $(C_{\bar{q}},s_{\bar{q}}^{i})$, as derived from Eq. (9), are due to the axial coupling $g_{A}^{q}$ of the $Z$-$q$-$\bar{q}$} interaction, which breaks charge symmetry and induces a forward-backward asymmetry, they vanish under the integration of $y$, and hence $\left\langle C_{q}\right\rangle=\left\langle C_{\bar{q}}\right\rangle$ and $\left\langle S_{q}^{i}\right\rangle=\left\langle S_{\bar{q}}^{i}\right\rangle$. This, together with Eq. (10), allows us to write Eq. (3) in a concise form,
@@
That leveraging the associatively produced hadron $h^{\prime}$ in a multichannel analysis enables one to disentangle the quark flavors of the dipole couplings is the most important observation in this paper. Without measuring $h^{\prime}$, which amounts to removing the $\bar{z}$ dependence and setting the FF to unity in Eq. (3), the $B^{i}$ vanishes identically in Eq. (14). To restore the transverse spin effects in this case requires not to integrate over $y$ in Eq. (11), and to utilize the forward-backward asymmetry associated with the charge symmetry breaking effects. However, such observable still constrains only a linear combination of the up and down quark dipole couplings, similarly to Refs. [20, 22], but not the individual ones.
V Numerical results and discussion
To simplify the numerical analysis, we further integrate over $(z,\bar{z},M_{h})$ in Eq. (12), resulting in a single-differential $\phi_{R}$ distribution,
[…]
@@ V Numerical results and discussion
As discussed, in order to separate the photon and $Z$ dipole couplings, we propose to carry out the measurements both at low-energy lepton colliders, such as the Super Tau-Charm Facility (STCF) [53], BELLE [54], and high-energy $Z$ factories, such as the Circular Electron-Position Collider (CEPC) [55, 56], International Linear Collider (ILC) [57], and the Future Circular Collider (FCC-ee) [58]. As a benchmark, we choose $\sqrt{s}=10\,{\rm GeV}$ with a polarized electron beam, with $\lambda_{\ell}=0.8$, to enhance the signal from the $\gamma\gamma$ channel to probe $\Gamma^{u,d}_{\gamma}$, and $\sqrt{s}=91\,{\rm GeV}$, at the $Z$-pole, with unpolarized lepton beams to probe the $\Gamma^{u,d}_{Z}$ from the $ZZ$ channel, both with an integrated luminosity of $\mathcal{L}=1000~{\rm fb}^{-1}$. We select for both cases the optimal kinematic regions, $z\in(0.4,0.8)$, $\bar{z}\in(0.2,0.8)$, and $M_{h}\in(0.4,1.4)~{\rm GeV}$, which minimize the uncertainties of the DiFFs in the global-fit analyses [45, 48, 41, 42, 43] and satisfy the factorization condition in Eq. (3). Within these regions, we estimate that the relative uncertainties from DiFFs in the spin asymmetries are at the level of less than 20%, with statistical errors being the dominant contribution. Consequently, the DiFF-related uncertainties have negligible impact on our conclusions. These uncertainties will be further reduced with inclusion of future data from the EIC and lepton colliders [59].
@@
The projected constraints on the light-quark dipole couplings $(\Gamma_{i}^{u},\Gamma_{i}^{d})$ are estimated from the asymmetries in Eq. (16) through a $\chi^{2}$ analysis, where the uncertainties are dominated by the statistical errors from Eq. (17). The resulting 68% confidence level limits are shown in Fig. 3 for (a) $i=\gamma$ and (b) $i=Z$, separately for the $\pi^{\pm}$ (red bands), $K^{\pm}$ (blue bands), and $p/\bar{p}$ (green bands) associated production channels. The real and imaginary parts are constrained in the same way. As expected, the linear dependence of $A_{LR,UD}^{h^{\prime}}$ on the dipole couplings leads to straight-band confidence regions; the $K^{\pm}$ channel only probes the $\Gamma_{\gamma,Z}^{u}$; and the $\pi^{\pm}$ and $p/\bar{p}$ give opposite signs for the slopes of the constraint bands. These multiple channels significantly constrain the dipole couplings, yielding a typical upper limit of $\mathcal{O}(10^{-2})$ for $\Gamma_{\gamma}^{u,d}$ and $\mathcal{O}(10^{-3})$ for $\Gamma_{Z}^{u,d}$, simultaneously for their real and imaginary parts. This complements the constraints from an $ep$ collider, especially when comparing with their flat directions of the constraining contour in the $(\Gamma_{i}^{u},\Gamma_{i}^{d})$ space [20, 22].
@@
While the dipole couplings $\Gamma_{\gamma,Z}^{u,d}$ can also be constrained by the Drell-Yan process, Higgs boson production via vector boson fusion, diboson production, associated production of weak and Higgs bosons at the LHC, low-energy data, and electroweak precision observables, the dipole couplings are only constrained quadratically at $\mathcal{O}(1/\Lambda^{4})$ when the transverse spin information associated with the initial or final state of the hard scattering processes is not considered. As a result, the sensitivity of those methods, among which Drell-Yan data provide the strongest constraint, is weaker than our results by one to two orders of magnitude, even with only one operator considered at a time [12, 13, 14, 15]. Furthermore, these methods lack the sensitivity to distinguish between the real and imaginary parts of dipole couplings and cannot separate them from other SMEFT operators. This highlights the advantage of the transverse spin observable proposed in this paper, which depends linearly and exclusively on the SMEFT dipole operators.
VI Conclusion
@@
In this paper, we introduced a novel approach to probe the light-quark dipole interactions through the azimuthal asymmetry between a collinear pair of hadrons produced in association with a well-separated third hadron at $e^{-}e^{+}$ colliders. This asymmetry results from the transverse spin of a final-state light quark, and is exclusively sensitive to the interference of the dipole and the SM interactions at the leading power of $\mathcal{O}(1/\Lambda^{2})$, unaffected by other new physics interactions in the hard scattering. This approach can simultaneously constrain the real and imaginary parts of the dipole couplings, provide a new way to probe $CP$ violation effects at $e^{-}e^{+}$ colliders, and outperform conventional methods proposed at the LHC and other observables at $e^{-}e^{+}$ colliders, by one to two orders of magnitude. By leveraging a hadron $h^{\prime}$ that is produced in association with the dihadron pair and combining all available channels of $h^{\prime}$ ($\pi^{\pm}$, $K^{\pm}$, and $p/\bar{p}$), we demonstrated—for the first time—that an $e^{-}e^{+}$ collider can resolve the flavor degeneracy present at $ep$ colliders, enabling separate constraints on the dipole couplings of up and down quarks. These developments open new avenues for probing NP at the precision frontier and establish an interdisciplinary connection between hadron structure studies and electroweak precision physics.
VII Acknowledgments